OpenStudy (anonymous):

Find the slope of (4,5) and (-2,5)

OpenStudy (mertsj):

\[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (anonymous):

i got 0/-6

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

you're close... not quite, but very close :)

OpenStudy (anonymous):

5-5=0 -2-4=-6

OpenStudy (anonymous):

*not quite right sorry forgot a word lol

OpenStudy (anonymous):

it is rise over run. So the two x's are run because they are horizontal so that is 6 because the different is six and the rise is 0 because the difference between the two y's is 0. So it is 0/6

OpenStudy (anonymous):

so simply it is just 0

OpenStudy (anonymous):

So its just 0

OpenStudy (anonymous):

 does that make sense?